An octahedron was used as dice in a special edition of snakes and ladders. A player throws the die twice and takes the sum before making a move. Find the mean and variance of the sum of two throws of the die.


1
Expert's answer
2021-05-10T18:29:58-0400

The distribution table for the sum of two dice is-





Mean "E(X) =\sum XP(X)=6.902"


Variance "= E(X^2)-[E(X)]^2"

"=54.074-(6.902)^2=54.704-47.637=7.06"